1. Find all square roots of the number. 256.
2. Use a calculator to approximate the square root. Round to three decimal places. -√0.00421.

Answer by CharlesG2(834)

1. Find all square roots of the number. 256.
2. Use a calculator to approximate the square root. Round to three decimal places. -√0.00421.
1. x^2 = 256
x = +- 16
2. x^2 = -sqrt(0.00421) (hoping this is what you asking for here)
x^2 = -1 * sqrt(0.00421)
x = +- 0.255i